1. Write the oxidation number for each atom: a. CrO42- b. HC104 d. Sg c. NF3 e. P2O5 g. CHC13

Write the oxidation number for each atom

Please help me solve these 6. I don't know how to do them.
Answer #1

In generally i) The sum of the oxidation states of all the atoms of ions in a neutral compound is equal to zero. The sum of@ crata Applicable dis rule and In generally oxidation number=-=; :: cF+ 4(-2) = -2 cv = 8-2 = +6 --+6 and 05-21 ordation n© NE In generally F = -1 N+36-1) = 0 N = +3 + N F + 3) -1 cidation numbers @ So 5o = s oxidation number is z ero. This is eA CHCl In generally cha-1 H +4 NÓ . C+1+36-1) =0 = c = + oxidation numbers H₃ +1 a -1 Cyta